# SAVeD: Learning to Denoise Low-SNR Video for Improved Downstream Performance

### v1

@@ B.1 SAVeD Architecture Details
Our method uses a series of convolution blocks with skip connections as an encoder $\Phi$, a bottleneck (hourglass network) $\Theta$, and a reconstruction decoder $\Psi$. Architectural details about each of these are shown in Tab. 6. For more implementation details, the code will be made publicly available.
@@
Table 6: Architecture details of the encoder, bottleneck, and decoder of SAVeD.Conv_block" is a basic convolutional block composed of 3x3 convolution with padding side of 1 and ReLU activation. “Skip" is a skip connection (stored to be input into the decoder) composed by maxpooling and then running a 1x1 convolution. “Upsample_block" is a 2D ConvTranspose with a 2x2 kernel and a stride of 2 and a ReLU activation. “Skip_connect” is the concatenation of the output from Upsample_block+Conv_block and the “Skip” corresponding to the same layer saved by the encoder. Note that this architecture is on input size of 1024x512.
@@
Table 7: Architecture details of the vanilla autoencoder.Conv_block" is a basic convolutional block composed of 3x3 convolution with padding side of 1 and ReLU activation. “Bilinear_upsample_block" is a Bilinear Upsample kernel with a scale factor of 2 and align corners set to True. Note that this architecture is on input size of 1024x512.
B.2 SAVeD Hyperparameters and Time Comparisons
The hyperparameters for our method are in Tab. 8. All DAE models are trained until the training loss converges on 2 NVIDIA RTX 4090 GPUs. Tab. 4 shows how much time each denoiser took on each dataset.
@@ B.2 SAVeD Hyperparameters and Time Comparisons
Table 8: SAVeD Hyperparameters. Note “$\text{inverse}(S_{t,T=1})$$=\min(0,I_{t}-I_{t-T})+I_{t}+\min(0,I_{t}-I_{t+T})$. f=Factor, pat=Patience, ss=Step size.
Figure 11: Visualization of $PSNR_{D_{KL}}$. Both images on the left are noisy images. The image on the far left has a fish located in the red bounding box. The image in the middle is a frame from the same video clip but with no fish in the red box. The histogram compares the pixel intensity values of the pixels within the bounding boxes. We can see these distributions, while overlapping, are distinct.
@@ B.3 CFC22 Detector Details
We fine-tune a YoloV5-small model pretrained on COCO using the default training settings from Ultralytics over 5 epochs with a batch size of 16. As in Kay et al. [37], we resize all inputs to have 896 pixels as their longest side; the learning rate is 0.0025. We select the best model checkpoint based on validation mAP50. We train on two NVIDIA RTX A6000 GPUs. We recognize that the number of epochs (5) differs from the number of epochs in the original paper (150), and that is two-fold: 1.) CFC22++ Val and Test Performance after 5 epochs are $<1\%$ lower than Val and Test Performance after 150 epochs, therefore our denoised improvement beats the CFC22++ method also after CFC22++ is trained for 150 epochs while the detector model based on SAVeD frames is trained for 5 epochs; 2.) We wanted to show that a very simple detector could be used as a result of passing in denoised frames.

### v2

@@ B.1 SAVeD Architecture Details
Our method uses a series of convolution blocks with skip connections as an encoder $\Phi$, a bottleneck (hourglass network) $\Theta$, and a reconstruction decoder $\Psi$. Architectural details about each of these are shown in Tab. 6. For more implementation details, the code is publicly available here.
@@
Table 6: Architecture details of the encoder, bottleneck, and decoder of SAVeD. “Conv_block” is a basic convolutional block composed of 3x3 convolution with padding side of 1 and ReLU activation. “Skip” is a skip connection (stored to be input into the decoder) composed by maxpooling and then running a 1x1 convolution. “Upsample_block” is a 2D ConvTranspose with a 2x2 kernel and a stride of 2 and a ReLU activation. “Skip_connect” is the concatenation of the output from Upsample_block+Conv_block and the “Skip” corresponding to the same layer saved by the encoder. Note that this architecture is on input size of 1024x512.
@@
Table 7: Architecture details of the vanilla autoencoder. “Conv_block” is a basic convolutional block composed of 3x3 convolution with padding side of 1 and ReLU activation. “Bilinear_upsample_block” is a Bilinear Upsample kernel with a scale factor of 2 and align corners set to True. Note that this architecture is on input size of 1024x512.
B.2 SAVeD Hyperparameter Comparisons
The hyperparameters for our method are in Tab. 8. All DAE models are trained until the training loss converges on 2 NVIDIA RTX 4090 GPUs.
@@ B.2 SAVeD Hyperparameter Comparisons
Table 8: SAVeD Hyperparameters. Note $\text{inverse}(S_{t,T=1})$$\min(0,I_{t}-I_{t-T})+I_{t}+\min(0,I_{t}-I_{t+T})$. f=Factor, pat=Patience, ss=Step size.
Figure 12: Visualization of FBDFBD. Both images on the left are noisy images. The image on the far left has a fish located in the red bounding box. The image in the middle is a frame from the same video clip but with no fish in the red box. The histogram compares the pixel intensity values of the pixels within the bounding boxes. We can see these distributions, while overlapping, are distinct.
@@ B.3 CFC22 Detector Details
We fine-tune a YoloV5-small model pretrained on COCO using the default training settings from Ultralytics over 5 epochs with a batch size of 16. As in Kay et al. [24], we resize all inputs to have 896 pixels as their longest side; the learning rate is 0.0025. We select the best model checkpoint based on validation mAP50. We train on two NVIDIA RTX A6000 GPUs. We recognize that the number of epochs (5) differs from the number of epochs in the original paper (150), and that is intentional. The reasoning is two-fold: 1.) CFC22++ Val and Test Performance after 5 epochs are $1\%$ lower than Val and Test Performance after 150 epochs, therefore our denoised improvement beats the CFC22++ method also after CFC22++ is trained for 150 epochs while the detector model based on SAVeD frames is trained for 5 epochs; 2.) We wanted to show that a very simple detector could be used as a result of passing in denoised frames.
@@ Appendix C Visualizations
Figure 16: RAFT [50] on CFC22 imagery and bounding box masks. On the left, we can see the optical flow signal does not find the fish. When looking at the motion from the bounding-box mask of the fish (making the background movement stationary), the optical flow signal area is far greater than the actual area of the fish. On the right are frames (with fish and corresponding bounding-box masks), when there are 12 fish in the frame at once. Again, optical flow’s signal is weak with the fish movement compared to the background. With the mask movement, optical flow signals cluster in groups of masked fish, but individuals are difficult to distinguish.